OpenStudy (anonymous):

PLEASE HELP Given sin -3/5 and x is in quadrant 3, what is the value of tan (x/2)

OpenStudy (anonymous):

Use the formula for tan (1/2 x) thats in your textbook. The formula requires you to plug in a value for cos x. Weel, if x is a 3rd quadrant angle, then cos x must be negative, as cos is negative in the 3rd quadrant. So if sin x = -3/5, so cos x = -4/5 Now plug in -4/5 for cos x, and finish the problem. Note, the final answer will have a minus sign in front of the square root.